If x=3 and x=-5, then form a quadratic equation.
Answer by stanbon(48558)

If x=3 and x=-5, then form a quadratic equation.
x-3=0 and x+5=0
Multiply to get:
(x-3)(x+5)=0
x^2+2x-15=0
You could write this as y=x^2+2x-15
Cheers,
Stan H.